# On spectrum of corona product of duplication signed graph and its application

### v4

@@ Abstract
This paper introduces the concept of $\mu$-signed and duplication signed graphs and shows that both are always structurally balanced. Using the duplication signed graph, we define the corona product of the duplication signed graph (Duplication add vertex corona product and duplication vertex corona product) and explore their structural properties. Additionally, we provide the adjacency spectrum of both the products for any $\Gamma_{1}$ and $\Gamma_{2}$, and the Laplacian and signless Laplacian spectrum for regular $\Gamma_{1}$ and arbitrary $\Gamma_{2}$, in terms of the corresponding spectrum of $\Gamma_{1}$ and $\Gamma_{2}$. Finally, we discuss its application in generating integral and equienergetic signed graphs.
@@ 1 Introduction
All graphs in this paper are finite, simple, and undirected. Let $\Gamma=(G,\sigma)$ be a signed graph with a vertex set $V=\{u_{1},u_{2},\ldots,u_{n}\}$ and an edge set $E=\{e_{1},e_{2},\ldots,e_{m}\}$, where $G=(V,E)$ is the underlying unsigned graph, $\sigma:E\rightarrow\{+,-\}$ is a mapping that assigns either a positive or a negative sign to the edges of $\Gamma$, known as the signature of $\Gamma$, and $n$ and $m$ are the order and size of $\Gamma$, respectively. The signed degree of vertex $u_{i}$ is denoted by $sdeg(u_{i})$, and it is given by $d_{u_{i}}^{+}-d_{u_{i}}^{-}$, where $d_{u_{i}}^{+}$ and $d_{u_{i}}^{-}$ are the number of positive and negative edges incident to $u_{i}$ respectively. The degree of vertex $u_{i}$ is given by $d_{u_{i}}=d_{u_{i}}^{+}+d_{u_{i}}^{-}$. A signed graph $\Gamma$ is called net-regular with net degree $k(k\in\mathbb{Z})$ if the signed degree of all the vertices of $\Gamma$ are $k$ [1], in addition $\Gamma$ is called co-regular if the underlying unsigned graph $G$ is r-regular and the pair $(r,k)$ is called the co-regularity pair [2].
A marking is a function $\mu:V\rightarrow\{1,-1\}$ which assigns either $+1$ or $-1$ to the vertices of $\Gamma$ and thus a signed graph is a $3$-tuple $\Gamma=(G,\sigma,\mu)$. Throughout this paper we will be discussing canonical marking, although multiple marking function of signed graphs exists [3]. For canonical marking $\mu(u_{i})=\prod_{e\in E_{u_{i}}}\sigma(e)$, where $E_{u_{i}}$ is the set of edges adjacent to $u_{i}$. Let $\Gamma_{1}=(\mathcal{G}_{1},\sigma_{1},\mu_{1})$ and $\Gamma_{2}=(\mathcal{G}_{2},\sigma_{2},\mu_{2})$ be two signed graphs. If $V(\mathcal{G}_{1})=\{u_{1},u_{2},\ldots,u_{n_{1}}\}$ and $V(\mathcal{G}_{2})=\{v_{1},v_{2},\ldots,v_{n_{2}}\}$ then the marking vector of $\Gamma_{1}$ and $\Gamma_{2}$ are column vectors of dimension $n_{1}\times 1$ and $n_{2}\times 1$ given by $\mu(\Gamma_{1})=[\mu_{1}(u_{1}),\mu_{1}(u_{2}),\ldots,\mu_{1}(u_{n_{1}})]^{T}$ and $\mu(\Gamma_{2})=[\mu_{2}(v_{1}),\mu_{2}(v_{2}),\ldots,\mu_{2}(v_{n_{2}})]^{T}$ respectively, where $A^{T}$ denotes transpose of $A$. Also we consider $\phi(\Gamma_{j})=diag[\mu(\Gamma_{j})]^{T}$ so that $\phi(\Gamma_{j})^{2}=I_{n_{j}}$ for $i=1,2.$
A cycle (C) is positive if the product of the signatures of its edges is positive. Otherwise, it is negative. A graph is balanced if all its cycles are positive [4]. The adjacency matrix of $\Gamma$, denoted by $A(\Gamma)=(a^{\sigma}_{ij})_{n\times n}$, is an $n\times n$ matrix, where $a_{ij}^{\sigma}=\sigma(u_{i}u_{j})a_{ij}$ and $a_{ij}=1$ if $u_{i}$ is adjacent to $u_{j}$, and 00 otherwise. The Laplacian matrix $L(\Gamma)$ and the signless Laplacian matrix $Q(\Gamma)$ are given by $L(\Gamma)=D(\Gamma)-A(\Gamma)$ and $Q(\Gamma)=D(\Gamma)+A(\Gamma)$, where $D(\Gamma)=diag(d_{u_{1}},d_{u_{2}},\ldots,d_{u_{n}})$, a diagonal matrix with entries $d_{u_{i}}$. Given the same underlying graph $\mathcal{G}$, two signed graphs, $\Gamma_{1}=(\mathcal{G},\sigma_{1})$ and $\Gamma_{2}=(\mathcal{G},\sigma_{2})$, are switching equivalent (referred to as $\Gamma_{1}\sim\Gamma_{2}$) if there exists a switching function $\theta$ such that for each edge $e=(u,v)$ in $\mathcal{G}$, $\theta(u)\sigma_{1}(e)\theta(v)=\sigma_{2}(e)$.
Two signed graphs are called $M$-cospectral if they have same $M$-spectrum, $M\in\{A(\Gamma),L(\Gamma),$
$Q(\Gamma)\}.$ The sum of absolute eigenvalues of $A(\Gamma)$ is called the energy of $\Gamma$ [5] and is denoted by
$E_{\Gamma}=\sum_{i=1}^{n}|\lambda_{i}|,$
where $\{\lambda_{i}:i=1,2,\ldots,n\}$ denotes the the eigenvalues of $A(\Gamma)$.
@@
In 1970 R Frucht and F Harary [6] first defined the corona product for unsigned graphs; later Cam McLeman and Erin McNicholas [7] first introduced the coronal of graphs to obtain the adjacency spectra for arbitrary graphs. Then Shu and Gui [8] generalized it and defined the coronal of the Laplacian and the signless Laplacian matrix of unsigned graphs. R.P. Varghese and D. Susha [9] introduced the duplication add vertex corona and duplication vertex corona spectra for unsigned graphs. We have built upon their research by exploring these spectra for signed graphs, and discovered that the two products are isomorphic. Additionally, we have analyzed the structural properties of the product using the framework established by Adhikari et al. [3]. Finally, we have also discussed the potential applications of these findings in generating equienergetic and integral signed graphs.
@@ 2 Preliminaries
Duplication graph of unsigned graph was defined by E Sampathkumar [10], following that we define the duplication signed graph as follows;
@@
Suppose $\Gamma(G,\sigma,\mu)$ be a signed graph with vertex set $V(\Gamma)=\{u_{1},u_{2},\ldots,u_{n}\}$ and $U(\Gamma)=\{a_{1},a_{2},\ldots,a_{n}\}$ be another set corresponding to $V(\Gamma)$ and having the same marking(i.e., $\mu(a_{i})=\mu(u_{i});\leavevmode\nobreak\ i=1,2,\ldots,n$). Draw $a_{i}$ adjacent to all the vertices in $N(u_{i})$, in $\Gamma$ (sign of new edge $e=\mu(a_{i})\mu(u),\leavevmode\nobreak\ u\in N(u_{i})$) and delete the edges of $\Gamma$ only. The signed graph thus obtained is called the duplication signed graph of $\Gamma$ and we denote it by $D\Gamma$.
Following the notion of marked strong signed graph structured defined by E.Sampathkumar et. al. [11] we define $\mu$-signed graph as follows;
@@
Let $\Gamma=(G,\sigma,\mu)$ be as signed graph with marking $\mu$, then the $\mu-$signed graph $\Gamma_{\mu}=(G,\sigma_{\mu},\mu)$ is a signed graph with same underlying graph G and marking $\mu$ but the following signature($\sigma_{\mu}$);
@@
Duplication signed graph of any graph is balanced.
@@
Let $\Gamma=(G,\sigma,\mu)$ be a signed graph and $D\Gamma$ be its duplication signed graph, then the Laplacian matrix of $D\Gamma$ is,
@@
So, 00 is one eigenvalue of $L(D\Gamma)$ as $\Gamma_{\mu}$ is balanced.
@@
Properties:
  1. (i)
    Kronecker product is associative.
  2. (ii)
    $(A\otimes B)^{T}=A^{T}\otimes B^{T}$.
  3. (iii)
    $(A\otimes B)(C\otimes D)=AC\otimes BD$, given the product $AC$ and $BD$ exists.
  4. (iv)
    $(A\otimes B)^{-1}=A^{-1}\otimes B^{-1}$, for non-singular matrices $A$ and $B$.
  5. (v)
    If $A$ and $B$ are $n\times n$ and $m\times m$ matrices, then det$(A\times B)=(detA)^{m}(detB)^{n}$.
@@ 3 Corona product of duplication signed graph
Let $\Gamma_{1}$ and $\Gamma_{2}$ be two vertex disjoint signed graphs with $n_{1}$ and $n_{2}$ vertices respectively. Let $D\Gamma_{1}$ be the duplication signed graph of $\Gamma_{1}$ with vertex set $V(\Gamma_{1})\cup U(\Gamma_{1})$, where $V(\Gamma_{1})=\{u_{1},u_{2},\ldots,u_{n_{1}}\}$ and $U(\Gamma_{1})=\{a_{1},a_{2},\ldots,a_{n_{1}}\}$. Duplication add vertex corona, $\Gamma_{1}\circledast\Gamma_{2}$, is the graph obtained from $D\Gamma_{1}$ and $n_{1}$ copies of $\Gamma_{2}$ by making $a_{i}$ adjacent to every vertices in the $i^{th}$ copy of $\Gamma_{2}$ for $i=1,2,\ldots,n_{1}.$
@@
Let $\Gamma_{1}$ and $\Gamma_{2}$ be two vertex disjoint signed graphs with $n_{1}$ and $n_{2}$ vertices respectively. Let $D\Gamma_{1}$ be the duplication signed graph of $\Gamma_{1}$ with vertex set $V(\Gamma_{1})\cup U(\Gamma_{1})$, where $V(\Gamma_{1})=\{u_{1},u_{2},\ldots,u_{n_{1}}\}$ and $U(\Gamma_{1})=\{a_{1},a_{2},\ldots,a_{n_{1}}\}$. Duplication vertex corona, $\Gamma_{1}\circledcirc\Gamma_{2}$, is the graph obtained from $D\Gamma_{1}$ and $n_{1}$ copies of $\Gamma_{2}$ by making $v_{i}$ adjacent to every vertices in the $i^{th}$ copy of $\Gamma_{2}$ for $i=1,2,\ldots,n_{1}.$
@@
$\begin{split}&f(u_{i})=a_{i};\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,\ldots,n_{1}\\ &f(a_{i})=u_{i};\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,\ldots,n_{1}\\ &f(v_{k})=v_{k};\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,2,\ldots,n_{2}.\end{split}$
From Definition 2.1, for each $u_{i}$ there exists corresponding $a_{i}$ and vice versa, so the above function is well defined and bijective. Also, for edges $(u_{i},a_{j})$, $(a_{i},v_{k})$, and $(u_{i},v_{k})$ in $\Gamma_{1}\circledast\Gamma_{2}$ there exists edges $(a_{j},u_{i})$, $(u_{i},v_{k})$, and $(a_{i},v_{k})$ in $\Gamma_{1}\circledcirc\Gamma_{2}$, respectively. Hence, the underlying graphs are isomorphic.
@@
Following Theorem 3.1, all the results of the subsequent section is true for both the product.
Figure 1: $A=P_{2}\circledast C_{3}$ and $B=P_{2}\circledcirc C_{3}$. The normal line implies a positive edge and the broken line implies a negative edge.

### v5

# On the corona product of duplication signed graph and its application

@@ Abstract
This paper introduces the concepts of $\mu$-signed graphs and duplication-signed graphs and demonstrates that both types are always structurally balanced. Using the framework of duplication-signed graphs, we define two new corona products: the duplication add vertex corona product and the duplication vertex corona product, and also show that they are switching isomorphic. The structural properties of these products are explored in detail. Additionally, we compute the adjacency spectrum for both corona products for any pair of signed graphs, $\Gamma_{1}$ and $\Gamma_{2}$, and derive the Laplacian and signless Laplacian spectra for regular $\Gamma_{1}$ and arbitrary $\Gamma_{2}$, expressed in terms of the spectra of the constituent signed graphs.
@@ 1 Introduction
Signed graphs are an important generalization of classical graph theory, where the edges of a graph are assigned a positive or negative sign. They have been studied very widely due to their applications in modeling structural balance in social networks, circuit theory, and various optimization problems. In the study of signed graphs, their spectral properties also play an important role as tools for understanding the structure and combinatorial features of these graphs. Let $\Gamma=(G,\sigma)$ be a signed graph with a vertex set $V=\{u_{1},u_{2},\ldots,u_{n}\}$ and an edge set $E=\{e_{1},e_{2},\ldots,e_{m}\}$, where $G=(V,E)$ is the underlying unsigned graph, $\sigma:E\rightarrow\{+,-\}$ is a mapping that assigns either a positive or a negative sign to the edges of $\Gamma$, known as the signature of $\Gamma$, and $n$ and $m$ are the order and size of $\Gamma$, respectively. The signed degree of vertex $u_{i}$ is denoted by $sdeg(u_{i})$, and it is given by $d_{u_{i}}^{+}-d_{u_{i}}^{-}$, where $d_{u_{i}}^{+}$ and $d_{u_{i}}^{-}$ are the number of positive and negative edges incident to $u_{i}$ respectively. The degree of vertex $u_{i}$ is given by $d_{u_{i}}=d_{u_{i}}^{+}+d_{u_{i}}^{-}$. A signed graph $\Gamma$ is called net-regular with net degree $k(k\in\mathbb{Z})$ if the signed degree of all the vertices of $\Gamma$ are $k$ [12]. In addition, $\Gamma$ is called co-regular if the underlying unsigned graph $G$ is r-regular. The pair $(r,k)$ is called the co-regularity pair [16]. A marking is a function $\mu:V\rightarrow\{1,-1\}$ which assigns either +1+1 or 1-1 to the vertices of $\Gamma$ and thus a signed graph is a $3$-tuple $\Gamma=(G,\sigma,\mu)$. Throughout this paper, we will discuss canonical marking, although multiple marking functions of signed graphs exist [1]. For canonical marking $\mu(u_{i})=\prod_{e\in E_{u_{i}}}\sigma(e)$, where $E_{u_{i}}$ is the set of edges adjacent to $u_{i}$. Let $\Gamma_{1}=(\mathcal{G}_{1},\sigma_{1},\mu_{1})$ and $\Gamma_{2}=(\mathcal{G}_{2},\sigma_{2},\mu_{2})$ be two signed graphs. If $V(\mathcal{G}_{1})=\{u_{1},u_{2},\ldots,u_{n_{1}}\}$ and $V(\mathcal{G}_{2})=\{v_{1},v_{2},\ldots,v_{n_{2}}\}$ then the marking vector of $\Gamma_{1}$ and $\Gamma_{2}$ are column vectors of dimension $n_{1}\times 1$ and $n_{2}\times 1$ given by $\mu(\Gamma_{1})=[\mu_{1}(u_{1}),\mu_{1}(u_{2}),\ldots,\mu_{1}(u_{n_{1}})]^{T}$ and $\mu(\Gamma_{2})=[\mu_{2}(v_{1}),\mu_{2}(v_{2}),\ldots,\mu_{2}(v_{n_{2}})]^{T}$ respectively, where $A^{T}$ denotes transpose of $A$. Also we consider $\phi(\Gamma_{j})=diag[\mu(\Gamma_{j})]^{T}$ so that $\phi(\Gamma_{j})^{2}=I_{n_{j}}$ for $i=1,2.$ A cycle (C) is positive if the product of the signatures of its edges is positive. Otherwise, it is negative. A graph is balanced if all its cycles are positive [9]. The adjacency matrix of $\Gamma$, denoted by $A(\Gamma)=(a^{\sigma}_{ij})_{n\times n}$, is an $n\times n$ matrix, where $a_{ij}^{\sigma}=\sigma(u_{i}u_{j})a_{ij}$ and $a_{ij}=1$ if $u_{i}$ is adjacent to $u_{j}$, and 0 otherwise. The Laplacian matrix $L(\Gamma)$ and the signless Laplacian matrix $Q(\Gamma)$ are given by $L(\Gamma)=D(\Gamma)-A(\Gamma)$ and $Q(\Gamma)=D(\Gamma)+A(\Gamma)$, where $D(\Gamma)=diag(d_{u_{1}},d_{u_{2}},\ldots,d_{u_{n}})$, a diagonal matrix with entries $d_{u_{i}}$. Given the same underlying graph $\mathcal{G}$, two signed graphs, $\Gamma_{1}=(\mathcal{G},\sigma_{1})$ and $\Gamma_{2}=(\mathcal{G},\sigma_{2})$, are switching equivalent (referred to as $\Gamma_{1}\sim\Gamma_{2}$) if there exists a switching function $\theta$ such that for each edge $e=(u,v)$ in $\mathcal{G}$, $\theta(u)\sigma_{1}(e)\theta(v)=\sigma_{2}(e)$. Two signed graphs are called $M$-cospectral if they have same $M$-spectrum, $M\in\{A(\Gamma),L(\Gamma),Q(\Gamma)\}.$ The sum of absolute eigenvalues of $A(\Gamma)$ is called the energy of $\Gamma$ [4] and is denoted by $E_{\Gamma}=\sum_{i=1}^{n}|\lambda_{i}|,$ where $\{\lambda_{i}:i=1,2,\ldots,n\}$ denotes the the eigenvalues of $A(\Gamma)$.
@@
In 1970, R Frucht and F Harary [7] first defined the corona product for unsigned graphs; later, Cam McLeman and Erin McNicholas [11] introduced the coronal of unsigned graphs to obtain the adjacency spectra for arbitrary graphs. Then Shu and Gui [5] generalized it and defined the coronal of the Laplacian and the signless Laplacian matrix of unsigned graphs. R.P. Varghese and D. Susha [19] introduced the duplication add vertex corona and duplication vertex corona spectra for unsigned graphs. We have built upon their research by exploring these spectra for signed graphs and discovered that the two products are switching isomorphic. Additionally, we have analyzed the structural properties of the product using the framework established by Adhikari et al. [1]. Lastly, we have discussed potential applications of these results to generating equienergetic and integral signed graphs. This work expands the study of signed graphs by enhancing both structural and spectral dimensions. In doing so, new opportunities open for further research in more specialized graph classes and their applications.
@@ 2 Preliminaries
The duplication graph of the unsigned graph was defined by E Sampathkumar [14]; following that, we define the duplication signed graph as follows;
@@
Suppose $\Gamma(G,\sigma,\mu)$ be a signed graph with vertex set $V(\Gamma)=\{u_{1},u_{2},\ldots,u_{n}\}$ and $U(\Gamma)=\{a_{1},a_{2},\ldots,a_{n}\}$ be another set corresponding to $V(\Gamma)$ and having the same marking(i.e., $\mu(a_{i})=\mu(u_{i});\penalty 10000\ i=1,2,\ldots,n$). Draw $a_{i}$ adjacent to all the vertices in $N(u_{i})$, in $\Gamma$ (sign of new edge $e=\mu(a_{i})\mu(u),\penalty 10000\ u\in N(u_{i})$) and delete the edges of $\Gamma$ only. The signed graph thus obtained is called the duplication signed graph of $\Gamma$, and we denote it by $D\Gamma$.
Following the notion of marked strong signed graph-structured defined by E.Sampathkumar et al. [15], we define $\mu$-signed graph as follows;
@@
Let $\Gamma=(G,\sigma,\mu)$ be a signed graph with marking $\mu$, then the $\mu$-signed graph $\Gamma_{\mu}=(G,\sigma_{\mu},\mu)$ is a signed graph with same underlying graph G and marking $\mu$ but the following signature($\sigma_{\mu}$);
@@
The duplication signed graph of any graph is balanced.
@@
Let $\Gamma=(G,\sigma,\mu)$ be a signed graph, and $D\Gamma$ be its duplication signed graph, then the Laplacian matrix of $D\Gamma$ is,
@@
So, 0 is one eigenvalue of $L(D\Gamma)$ as $\Gamma_{\mu}$ is balanced.
@@
The Kronecker product has several important properties:
  1. 1.
    It is associative: for any matrices $A,B,C,$ and $D$, we have $(A\otimes B)\otimes C=A\otimes(B\otimes C)$.
  2. 2.
    The transpose of the Kronecker product can be expressed as: $(A\otimes B)^{T}=A^{T}\otimes B^{T}.$
  3. 3.
    When multiplying Kronecker products, the following holds: $(A\otimes B)(C\otimes D)=AC\otimes BD,$ as long as the matrix products $AC$ and $BD$ are defined.
  4. 4.
    For non-singular matrices $A$ and $B$, the inverse of the Kronecker product is: $(A\otimes B)^{-1}=A^{-1}\otimes B^{-1}.$
  5. 5.
    If $A$ is an $n\times n$ matrix and $B$ is an $m\times m$ matrix, the determinant of the Kronecker product is given by: $\text{det}(A\otimes B)=(\text{det}A)^{m}(\text{det}B)^{n}.$
@@ 3 Corona product of duplication signed graph
Let $\Gamma_{1}$ and $\Gamma_{2}$ be two vertex-disjoint signed graphs with $n_{1}$ and $n_{2}$ vertices, respectively. Let $D\Gamma_{1}$ be the duplication signed graph of $\Gamma_{1}$ with vertex set $V(\Gamma_{1})\cup U(\Gamma_{1})$, where $V(\Gamma_{1})=\{u_{1},u_{2},\ldots,u_{n_{1}}\}$ and $U(\Gamma_{1})=\{a_{1},a_{2},\ldots,a_{n_{1}}\}$. Duplication add vertex corona, $\Gamma_{1}\circledast\Gamma_{2}$, is the graph obtained from $D\Gamma_{1}$ and $n_{1}$ copies of $\Gamma_{2}$ by making $a_{i}$ adjacent to every vertices in the $i^{th}$ copy of $\Gamma_{2}$ for $i=1,2,\ldots,n_{1}.$
@@
Let $\Gamma_{1}$ and $\Gamma_{2}$ be two vertex-disjoint signed graphs with $n_{1}$ and $n_{2}$ vertices, respectively. Let $D\Gamma_{1}$ be the duplication signed graph of $\Gamma_{1}$ with vertex set $V(\Gamma_{1})\cup U(\Gamma_{1})$, where $V(\Gamma_{1})=\{u_{1},u_{2},\ldots,u_{n_{1}}\}$ and $U(\Gamma_{1})=\{a_{1},a_{2},\ldots,a_{n_{1}}\}$. Duplication vertex corona, $\Gamma_{1}\circledcirc\Gamma_{2}$, is the graph obtained from $D\Gamma_{1}$ and $n_{1}$ copies of $\Gamma_{2}$ by making $v_{i}$ adjacent to every vertices in the $i^{th}$ copy of $\Gamma_{2}$ for $i=1,2,\ldots,n_{1}.$
@@
$\begin{split}&f(u_{i})=a_{i};\penalty 10000\ \penalty 10000\ for\penalty 10000\ \penalty 10000\ i=1,2,\ldots,n_{1}\\ &f(a_{i})=u_{i};\penalty 10000\ \penalty 10000\ for\penalty 10000\ \penalty 10000\ i=1,2,\ldots,n_{1}\\ &f(v_{k})=v_{k};\penalty 10000\ \penalty 10000\ for\penalty 10000\ \penalty 10000\ k=1,2,\ldots,n_{2}.\end{split}$
From Definition 2.1, for each $u_{i}$, there exists a corresponding $a_{i}$ and vice versa, so the above function is well-defined and bijective. Also, for edges $(u_{i},a_{j})$, $(a_{i},v_{k})$, and $(u_{i},v_{k})$ in $\Gamma_{1}\circledast\Gamma_{2}$ there exists edges $(a_{j},u_{i})$, $(u_{i},v_{k})$, and $(a_{i},v_{k})$ in $\Gamma_{1}\circledcirc\Gamma_{2}$, respectively. Hence, the underlying graphs are isomorphic.
@@
Following Theorem 3.1, all the results of the subsequent section are true for both products.
Figure 1: $A=P_{2}\circledast C_{3}$ and $B=P_{2}\circledcirc C_{3}$. The normal line implies a positive edge, and the broken line implies a negative edge.